Given n=1010 and p= 0.84, determine whether this is approxamlty a normal distrubution. if it is find the standard deviation of the sampling distribution of proportion.
Answer #1

Solution

Given that,

p = 0.84

1 - p =1-0.84=0.16

n = 1010

standard deviation =\sigma\hat p =  \sqrt[p ( 1 - p ) / n] =  \sqrt [(0.84*0.16) / 1010 ] = 0.0115

(*)when n = 10

than

standard deviation =\sigma\hat p =  \sqrt[p ( 1 - p ) / n] =  \sqrt [(0.84*0.16) / 10 ] = 0.1159